Company XYZ has a monopoly on a new product. The product can be manufactured in either of two factories. The costs of production for the two factories are MC1 = 10 + Qi and MC2 = 3 + 3Q2. The firm's demand for the product is P= 10 – 2(Q1 + Q2). How much should the firm produce in each factory? At what price should it sell the product?
Answer #1

Total reuomes TRF Marginal revues MR= MR at any point. Hice, on a new product, compay XYZ has a monopoly The films demand fothat, bactory 1 should produce o nit, factory 2 should produce, 1 nits To ful output 1 nt =Q. puce p= 10- 2Q2 = 10-2 =$8 = C